OpenStudy (anonymous):

Minimum Wage The table shows the minimum wage for three different years. Year/wages: 1940/0.25, 1968/1.60, 1997/5.15 (a) Make a scatterplot of the data in the viewing rectangle [1930, 2010, 10] by [0, 6, 1]. (b) Find a quadratic function given by that models the data. f(x)=a(x - h)^2 + k (c) Estimate the minimum wage in 1976 and compare it to the actual value of $2.30. (d) Estimate when the minimum wage was $1.00. (e) If current trends continue, predict the minimum wage in 2009. Compare it to the projected value of $7.25.

OpenStudy (anonymous):

can someone help me solve this so i can understand it

OpenStudy (amistre64):

it looks like they give you 3 points of reference, right?

OpenStudy (anonymous):

yea

OpenStudy (amistre64):

this turn into a system of 3 equations with 3 unknown variables; the unknowns are the constant coefficients of our quadratic equation that we are trying to find

OpenStudy (amistre64):

lets adjust the years so that the computations are smaller, but equivalent 1940/0.25, 1968/1.60, 1997/5.15 -1940 -1940 -1940 ------------------------------- 0, 0.25 28, 1.60 57, 5.15 in essense, all ive done by this is to shift the graph so to the left; same shape, just a different location 

OpenStudy (amistre64):

makes sense so far?

OpenStudy (anonymous):

yes so far

OpenStudy (amistre64):

i could also move the graph up or down if need be by adjusting the "y" values, but they look simple enough that i wont try that .... yet :)

OpenStudy (anonymous):

oh boy ok math isn't my strongest subject

OpenStudy (amistre64):

0, 0.25 28, 1.60 57, 5.15 when x=0, y = 0.25 ; this is our y intercept by definition, whenever x = 0 we have a y intercept ax^2 +bx + 0.25 = 0.25 ; when x = 0 ax^2 +bx + .25 = 1.60 ; when x = 28 ax^2 +bx + .25 = 5.15 ; when x = 57

OpenStudy (amistre64):

if math aint your strongest subject then i take it trying a matrix on this wont make alot of sense to you :)

OpenStudy (anonymous):

probably not..i can usally figure it out but for some reason i can't wrap my head around this problem i appreciate your help

OpenStudy (amistre64):

lets apply our x values in each case and see if we can solve for a and b 0a +0b + 0.25 = 0.25 ; this ones not going to give us any new information so we can ignore it for now 28^2a +28b = 1.60-.25 57^2a +57b = 5.15-.25 28^2a +28b = 1.35 57^2a +57b = 4.90 now we are left with a system of 2 equations in 2 unknowns to solve for a and b

OpenStudy (amistre64):

if we forget about trying to find the smallest value of this and that, we can simply eliminate some prospects by multiplying the top by 57^2 and the bottom by -28^2 28^2 (57^2)a +28 (57^2)b = 1.35(57^2) 57^2(-28^2)a +57(-28^2)b = 4.90(-28^2) --------------------------------------- 0 +(28 (57^2)+57(-28^2))b = 1.35(57^2) + 4.90(-28^2) dividing off that messy stuff gives us: \[b=\frac{1.35(57^2) + 4.90(-28^2)}{28 (57^2)+57(-28^2) }\]

OpenStudy (anonymous):

if the problem is ax^2+bx +.25 would the equation be a(28)^2+b(28)+.25? or am i wrong in assuming that?

OpenStudy (amistre64):

with the aid of a calculator; b = abt .012 and the same concept can be applied to get rid of the b to solve for a

OpenStudy (amistre64):

your right, I just subtracted the .25 from each side to get a and b all alone

OpenStudy (anonymous):

ok im confused 28^2 (57^2)a +28 (57^2)b = 1.35(57^2) 57^2(-28^2)a +57(-28^2)b = 4.90(-28^2) --------------------------------------- 0 +(28 (57^2)+57(-28^2))b = 1.35(57^2) + 4.90(-28^2) dividing off that messy stuff gives us: \[b=\frac{1.35(57^2) + 4.90(-28^2)}{28 (57^2)+57(-28^2) }\] this comes from ?

OpenStudy (amistre64):

28^2(57)a + 28(57)b = 1.35(57) 57^2(-28)a +57(-28)b = 4.90(-28) ------------------------------- (28^2(57)+57^2(-28))a +0 = 1.35(57) + 4.90(-28) \[a=\frac{1.35(57) + 4.90(-28)}{28^2(57)+57^2(-28)}=-60.25\] ill explain the best i can after i post this bit of it

OpenStudy (amistre64):

the numbers are messy, but the concept is still to use elimination to get rid of one of the variables in order to solve whats left for the remaing variable

OpenStudy (amistre64):

\[\begin{array}\ a_1x+b_1c=n\\a_2x+b_2c=m\end{array}\] \[\begin{array}\ (a_1x+b_1c=n)*a_2\\(a_2x+b_2c=m)*(-a_1)\end{array}\] \[\begin{array}\ \cancel{a_1a_2}x+a_2b_1c=a_2n\\\bcancel{-a_1a_2}x-a_1b_2c=-a_1m\\------------\\(a_2-a_1)b_2c=a_2n-a_1m\end{array}\]

OpenStudy (amistre64):

ugh, forgot a few parts in that; but it should still make sense i hope lol

OpenStudy (anonymous):

not really im an epic fail at this point of no sleep

OpenStudy (amistre64):

thats ok, these things can get complicated; i just tested out my "solution" and apparently i missed it someplace.

OpenStudy (anonymous):

ok well im lost so i cant help you much :/

OpenStudy (amistre64):

i hit a wrong button for the "a"; with any luck its about .001 so i want to test out the following to see how close I get to the points given y = .001x^2 +.012x+.25

OpenStudy (anonymous):

ok i'll be right here :)

OpenStudy (amistre64):

its yelling at me :/ with my approximations the wolf is giving me a value that is .25 cents less. ima gonna try the matrix route and see what it gives me .... but the concepts i presented are sound, its just working the numbers themselves thats a bit helter skelter

OpenStudy (anonymous):

oh don't let it get you like it has gotten me I still need to explain this to others I am working with so the better you are the more it helps me :) cuz right now i am a deer in head lights just looking at the equation coming

OpenStudy (amistre64):

which is what i had to begin with .... so i think the wolf is getting a little mixed up these days

OpenStudy (anonymous):

oh my

OpenStudy (anonymous):

:/

OpenStudy (amistre64):

the "exact" values for our coeefs are monster looking fractions; and my approximation to those is making my quadratic miss the mark the further it moves away from 1940

OpenStudy (anonymous):

:( oh no

OpenStudy (amistre64):

\[y = \frac{241}{185136}(x-1940)^2 +\frac{10891}{925680}(x-1940)+\frac{1}{4}\] that works lol

OpenStudy (anonymous):

omg what is that

OpenStudy (amistre64):

thats the quadratic equation that fits your data points; i got no idea how that would look in the a(x-h)^2+k format tho

OpenStudy (anonymous):

hmm

OpenStudy (amistre64):

for x=1997 i get 2.36; and the actual they presented in 2.30

OpenStudy (amistre64):

it estimates about 1960 for a minimum wage of 1.00

OpenStudy (amistre64):

and it estimates 7.26 for 2009

OpenStudy (anonymous):

ok well thats awesome thank you know i just have to figure out how I am going to show and explain it

OpenStudy (amistre64):

i think youd have to use computer generated values in order to "show" it; but the explanation is the basic notion that i started with

OpenStudy (amistre64):

we create a system of 3 equations in 3 unknown and solve it using approved methods

OpenStudy (amistre64):

either substitution, elimination, or matrix methods

OpenStudy (anonymous):

ok thank you at least you got me headed in the right direction :)

OpenStudy (amistre64):

lol, good luck with it ;)

OpenStudy (anonymous):

thanks